### v1

# Is there a dynamical tendency in $H_{0}$ with late time measurements?

Abstract
The discrepancy between the Hubble constant $H_{0}$ values derived from early-time and late-time measurements, reaching up to $4\sigma$, represents the most serious challenge in modern cosmology and astrophysics. In this work, we investigate if a similar tension exists between only late time measurements at different redshifts. We use the latest public datasets including Cosmic Chronometers, Megamasers, SNe Ia and DESI-BAO, that span from redshift $z\sim 0$ up to $z\sim 2.3$. By dividing the data into redshift bins, we derive $H_{0}$ values from each bin separately. Our analysis reveals a phenomenological dynamic evolution in $H_{0}$ across different redshift ranges, with a significance from $1.5\sigma$ and $2.3\sigma$, depending on the parameterization. Consistency of the model demands observational constancy of $H_{0}$ since it is an integration constant within the Friedmann-Lemaître-Robertson-Walker (FLRW) metric. Thus, these findings suggest that the observed Hubble tension might not only exist between early and late-time measurements but also among late-time data themselves, providing new insights into the nature of the Hubble tension.
1 Introduction
Our current understanding of the Universe relies on the $\Lambda$CDM model which incorporates the cosmological constant $\Lambda$ and a cold dark matter component. This model is the most accepted and successful in explaining most cosmological observations [1, 2]. One of the most important challenges within the astrophysics and cosmology community is the discrepancy of more than $4\sigma$ for the Hubble constant $H_{0}$, between local measurements using the cosmic distance ladder and early-time measurements of the cosmic microwave background (CMB). The former based on direct observations of Supernovae Type Ia (SNe Ia) calibrated with Cepheids by the SH0ES team [3] and the latter obtained by the Planck satellite, which relies on the $\Lambda$CDM model [1]. This so called Hubble tension has led to an exhaustive search for potential explanations. A wide range of models have been used to address the $H_{0}$ tension, proposing either a dark energy (DE) component during the early evolution of the Universe [4, 5, 6, 7, 8, 9], a DE component with a time-varying equation of state [10, 11, 12], extra interactions between the components of the Universe [13, 14, 15, 16], or modified gravity [17, 18, 19]. For a more comprehensive study about the $H_{0}$ tension and alternative proposals, see the reviews [20, 21, 22].
In the $\Lambda$CDM model, $H_{0}$ is a constant by definition: it represents the Hubble parameter $H(z)$ at the present epoch $H_{0}=H(z=0)$. However, recent results have shown that $H_{0}$ depends on the redshift of the probes used to determine it. The first of these results, from six gravitationally lensed quasars with measured time delays by the H0LiCOW program, indicated that $H_{0}$ decreases with lens redshift [23], with no evidence suggesting that this trend is due to unaccounted systematics [24]. Subsequent studies have found a similar feature in different datasets [25, 26, 27, 28, 29, 30, 31, 32, 33, 34], offering a new perspective on the Hubble tension due to an apparent redshift dependence of a quantity that should, tautologically, be a constant.
Recently, the Dark Energy Spectroscopic Instrument (DESI) collaboration published its cosmological constraints from the first year of observations of baryon acoustic oscillations (BAO). The results favor models with a DE component whose equation of state varies with time by more than $2.6\sigma$ [35], which challenges the paradigm of what we previously believed to be correct in the $\Lambda$CDM model. This dynamical DE could be responsible for a $H_{0}$ that depends on redshift.
Ref. [25] found that splitting cosmological data in different redshift bins up to $z=0.7$ gives different estimates to $H_{0}$ which are in contradiction with a single constant $H_{0}$ at $2.1\sigma$ level. Leading to the conclusion that the Hubble tension is present even between late time measurements themselves when taken at different redshifts. Ref. [25] assumed a cosmological constant and used BAO-BOSS data [2]. The new DESI results lead us to question whether the new data together with a dynamical DE could solve this tension.
The paper is organized as follows. In section 2, we detail the implemented methodology, along with the theory behind the cosmological probes used in this paper. In section 3, we mention the data used. In section 4, we explain the criteria used to split the data into bins. In section 5, we present our results on constraining the value of $H_{0}$ in each of the bins. Finally, the summary and conclusions can be found in section 6.
2 Cosmology
In light of the new discoveries for the first cosmological results from DESI [35], in this work we use a flat Universe with a dynamical Dark Energy component, with a Chevallier-Polarski-Linder (CPL) equation of state given by $w(a)=w_{0}+w_{a}(1-a)$ [36, 37]. The Friedmann equation is given by the following parameterization
$H(z)=H_{0}\left[\Omega_{m}(1+z)^{3}+(1-\Omega_{m})f(z)\right]^{1/2},$ (2.1)
where
$f(z)=(1+z)^{3(1+w_{0}+w_{a})}e^{-3w_{a}z/(1+z)},$ (2.2)
and with $H_{0}$, $\Omega_{m}$, $w_{0}$, $w_{a}$ the free parameters.
Following [25], we will split different cosmological data according to their redshift in order to estimate the Hubble constant using only data in certain redshift ranges. This method involves binning a modern dataset that includes the new DESI data, cosmic chronometers (CC), megamasers, and SNe Ia samples. The SNe Ia samples comprises the Pantheon+ compilation, and the new Union3 compilation and Dark Energy Survey (DES) Year 5 data release. With this new data we were able to expand the redshift range in the analysis to observe in more depth the evolution of $H_{0}$ with the redshift of the data.
Cosmic chronometers provide a direct estimate of the expansion rate of the Universe. This method takes advantage of the fact that the Hubble parameter, $H(z)$, can be directly expressed as a function of the time differential of the Universe, $dt$, over a given redshift interval, $dz$ [38]
$H(z)=-\frac{1}{1+z}\frac{dz}{dt}.$ (2.3)
CCs evolve homogeneously as a function of cosmic time and on a timescale much longer than their age difference. Given $H(z)$, to obtain $H_{0}$ from eq. (2.1) it is necessary to determine $\Omega_{m}$, $w_{0}$, $w_{a}$ from other observations.
The water megamasers found in the accretion disks of supermassive black holes in active galactic nuclei (AGN) act like a laser beam in the microwave band, providing a method to calculate extragalactic distances without relying on the cosmic distance ladder [39]. The disk model returns an estimation of the angular diameter distance $\hat{D}_{\text{A}}$ to the galaxy and its recession velocity $\hat{v}_{i}$. The decomposition of this velocity between its peculiar and cosmological components $\hat{v}_{i}=v_{\textrm{pec,i}}+cz_{i}$ is unknown for objects this close. Each $\hat{D}_{\text{A}}$ is related to its expected redshift $z_{i}$ and the cosmological parameters by the equation
$D_{\textrm{A}}(z)=\frac{c}{1+z}\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}.$ (2.4)
Following [40], we take peculiar velocities of the galaxies into account incorporating $\sigma_{\textrm{pec}}=250$ km $\textrm{s}^{-1}$ into the velocity uncertainties. The chi-squared used to constrain cosmological parameters is then given as
$\chi^{2}=\sum^{N}_{i=1}\left[\frac{(v_{i}-\hat{v}_{i})^{2}}{\sigma_{v,i}^{2}+\sigma_{\textrm{pec}}^{2}}+\frac{(D_{\textrm{A}}(v_{i}/c)-\hat{D}_{i})^{2}}{\sigma_{D,i}^{2}}\right],$ (2.5)
where the expected cosmological recession velocities are given by $v_{i}=cz_{i}$, $N$ the number of data, $\sigma_{v,i}$ is the statistical uncertainty in the velocity measurement $\hat{v}_{i}$, and $\sigma_{D,i}$ is the standard deviation of the distance measurement $\hat{D}_{i}$. We take the expected velocities $v_{i}$ for megamaser host galaxies as nuisance parameters.
The SNe Ia data gives the apparent magnitude $m_{\textrm{b}}$ of each SNe, which can be compared with the predicted model given by
$m_{\textrm{b}}=\textrm{M}+\mu(z)=\textrm{M}+25+5\textrm{log}_{10}\left(\frac{D_{\textrm{L}}(z)}{\textrm{Mpc}}\right),$ (2.6)
where M is the absolute magnitude for SNe Ia, and $D_{\textrm{L}}(z)$ is the luminosity distance, which in the case of SNe Ia can be written, according to [41], by
$D_{\textrm{L}}(z)=(1+z_{\textrm{hel}})(1+z_{\textrm{cmb}})D_{\textrm{A}}(z_{\textrm{cmb}}),$ (2.7)
$z_{\textrm{hel}}$ is the heliocentric redshift of the SNe (or its host galaxy) and $z_{\textrm{cmb}}$ is this redshift corrected for the peculiar motion of the solar system with respect to the CMB. Note that the parameters M and $H_{0}$ are degenerate when analyzing SNe alone. Therefore, SNe measurements are used to constrain other cosmological parameters, particularly the matter density $\Omega_{m}$.
The BAO measurements provide distances relative to the sound horizon, $r_{d}$, as a function of redshift [2]. Along the line-of-sight direction, a measurement of the redshift interval, $\Delta z$, gives a means to measure the Hubble distance at redshift $z$,
$D_{\textrm{H}}(z)=\frac{c}{H(z)}.$ (2.8)
Along the transverse direction, measuring the angle $\Delta\theta$ subtended by the BAO feature at $z$ estimates the comoving angular diameter distance, $D_{\text{M}}(z)$
$D_{\textrm{M}}(z)=(1+z)D_{\textrm{A}}(z),$ (2.9)
These measurements can also be summarized by a single estimation representing the spherically averaged distance
$D_{\textrm{V}}(z)=\left[zD^{2}_{\textrm{M}}(z)D_{\textrm{H}}(z)\right]^{1/3},$ (2.10)
BAO measurements constrain these distances all divided by the sound horizon $r_{\textrm{d}}$. Therefore, $H_{0}$ and $r_{\textrm{d}}$ are degenerated with each other. The BAO measurements constraint the product $H_{0}r_{\textrm{d}}$, and therefore are key in the treatment of Hubble tension. This constraint implies that a higher $H_{0}$, as that obtained by SH0ES [3], requires a lower $r_{\text{d}}$, which is incompatible with the value inferred by the Planck Collaboration [1]. In this work we won’t use the CMB prior on $r_{\textrm{d}}$ as this will imply the use of data outside our redshift bins.
3 Data
To constrain $H_{0}$, we use the following public data up to the redshift $z<2.33$, including their systematic errors and covariance matrices. These data correspond to the latest public measurements of the observables from the previous section:
  1. I.
    Distances and recession velocities from 6 megamasers in the range $0.002\leq z\leq 0.034$, from the Megamaser Cosmology Project [39, 40, 42].
  2. II.
    32 cosmic chronometers data in the range $0.07\leq z\leq 2$, from refs. [43, 44, 45, 46, 47, 48, 49, 50, 51]. We estimate the covariance matrix [^1] The detailed recipe on how to correctly estimate the covariance matrix for cosmic chronometers is given in https://gitlab.com/mmoresco/CCcovariance between these data points following ref. [52].
  3. III.
    BAO measurements in galaxy, quasar and Lyman-$\alpha$ forest tracers from the Year 1 data release of DESI. These are measurements of the transverse comoving distance $D_{\textrm{M}}$ and Hubble rate $D_{\textrm{H}}$, or their combination $D_{\textrm{V}}$, relative to the sound horizon $r_{\textrm{d}}$ with the following effective redshifts $z_{\textrm{eff}}=0.30,0.51,0.71,0.93,1.32,1.49,2.33$ [35].
  4. IV.
    For the superenovae type Ia measurements, we use any of the three latest compilations. 1,590 SNe Ia from the Pantheon+ compilation in the redshift range $0.01\leq z\leq 2.26$ [53]. Recently published results from the Year 5 data release of the Dark Energy Survey (DES Y5) that include 1,635 new photometrically classified SNe Ia and 194 low-redshift SNe Ia (which are also in common with Pantheon+), spanning $0.025\leq z\leq 1.3$ [54]. Finally, 22 binned data points from 2087 SNe Ia from the Union3 sample in the range $0.05\leq z\leq 2.26$ [55], which has 1360 SNe in common with Pantheon+. We include both the statistical and systematic uncertainties in the three sets. We emphasize that these SNe Ia data are not independent of each other, and therefore, we cannot combine them. However, they differ in the way they analyze systematic errors and astrophysical parameters, making it interesting to compare the results of each of them.
Compared with the data used in ref. [25] which includes the BOSS BAO data [2] and the Pantheon sample [56], we have the BAO data from DESI Y1, the recent supernovae compilations Pantheon+, Union3 and DES Y5, as well as three extra measurements of CC.
Since the Pantheon+, DES and Union3 data cannot be combined, we present each of their results separately. The analyses always include the megamasers+CC+DESI data, which we will call “Base”, but differ in the SNe sample. We denote each case as Base+PantheonPlus, Base+Union3 and Base+DES.
4 Fitting
Base+PantheonPlus
Bin Data Range $\bar{z}$
1 Megamasers, SNe $0.01<z\leq 0.069$ $0.032$
2 SNe, CC $0.069<z\leq 0.199$ $0.15$
3 SNe, CC, DESI-BGS $0.199<z\leq 0.425$ $0.30$
4 SNe, CC, DESI-LRG1 ($z=0.51$) $0.425<z\leq 0.625$ $0.51$
5 SNe, CC, DESI-LRG2 ($z=0.71$) $0.625<z\leq 0.7891$ $0.71$
6 SNe, CC, DESI-LRG3+ELG1 $0.7891<z\leq 1.13$ $0.93$
7 SNe, CC, DESI-ELG2, DESI-QSO $1.13<z\leq 1.65$ $1.38$
8 SNe, CC, DESI-Lya $1.65<z\leq 2.3$ $1.94$
Base+Union3
Bin Data Range $\bar{z}$
1 Megamasers, SNe $0<z\leq 0.069$ $0.05$
2 SNe, CC $0.069<z\leq 0.19$ $0.12$
3 SNe, CC, DESI-BGS $0.19<z\leq 0.44$ $0.30$
4 SNe, CC, DESI-LRG1 ($z=0.51$) $0.44<z\leq 0.649$ $0.51$
5 SNe, CC, DESI-LRG2 ($z=0.71$) $0.649<z\leq 0.79$ $0.71$
6 SNe, CC, DESI-LRG3+ELG1 $0.79<z\leq 1.23$ $0.93$
7 SNe, CC, DESI-ELG2, DESI-QSO $1.23<z\leq 1.7$ $1.36$
8 SNe, CC, DESI-Lya $1.7<z\leq 2.3$ $2.26$
Base+DES
Bin Data Range $\bar{z}$
1 Megamasers, SNe $0<z\leq 0.069$ $0.041$
2 SNe, CC $0.069<z\leq 0.19$ $0.12$
3 SNe, CC, DESI-BGS $0.19<z\leq 0.401$ $0.30$
4 SNe, CC, DESI-LRG1 ($z=0.51$) $0.401<z\leq 0.63$ $0.51$
5 SNe, CC, DESI-LRG2 ($z=0.71$) $0.63<z\leq 0.826$ $0.71$
6 SNe, CC, DESI-LRG3+ELG1 $0.826<z\leq 1.13$ $0.93$
Table 1: Summary of the data, range in redshift and $\bar{z}$ in each bin for each different data set.
In this section, we describe how we fit the cosmological parameters using the data specified in the last section. Our objective is to see if the estimation of the Hubble constant $H_{0}$ is consistent when using data at different redshifts. We will split the total redshift range from our data in separate redshift bins with $z\in(z^{i}_{min},z^{i}_{max})$ and we use only data coming from that bin to estimate $H_{0}$.
The redshift width of the bins varies between each other as we follow the strategy used in [25] requiring that the average redshift of each bin coincides between megamasers, SNe and BAO. We use the weighted average $\bar{z}$ given by
$\bar{z_{i}}=\frac{\sum_{k}^{N_{i}}z_{k}(\sigma_{k})^{-2}}{\sum_{k}^{N_{i}}(\sigma_{k})^{-2}}\,,$ (4.1)
where $N_{i}$ denotes the number of data used in the bin $i$ and $\sigma_{k}$ denotes the error in the observable at redshift $z_{k}$. For megamasers we use $\sigma_{k}^{2}=\sigma_{\hat{v},k}^{2}+\sigma_{\textrm{pec}}^{2}+\sigma_{\hat{D},k}^{2}$. The bins constructed in each analysis are summarized in table 1.
Not all bins contain all the data. The megamasers appear only in the first redshift bin, while some bins lack DESI data. We only require that all bins contain either CCs or megamasers to break the degeneracies needed to obtain a value of $H_{0}$. We emphasize that the condition (4.1) leads to different binnings and different $\bar{z}$’s depending on the SNe’s used (PantheonPlus, Union3 and DES).
To fit the parameters, we used the Python module emcee [57]. We set flat priors for all the parameters, in particular for $H_{0}$ the prior is large enough that its fitting is mainly influenced by the measurements
$H_{0}\in(0,150).$ (4.2)
On the other hand, the data contained in a single bin lack the power to constrain the equation of state parameters $w_{0}$ and $w_{a}$. Therefore the posteriors of these parameters are mainly influenced by their priors. We considered large prior ranges
$w_{a}\in(-2.5,2.5),\;\;w_{0}\in(-1.5,-1/3),$ (4.3)
to allow for enough freedom in the DE dynamics. These ranges contain both the DESI estimation [^2] We recall that $w_{0}=-0.64\pm 0.11$, and $w_{a}=-1.27^{+0.40}_{-0.34}$ for DESI+CMB+Union3 [35]. and the cosmological constant $w_{0}=-1$, $w_{a}=0$. After we fit the parameters, $w_{0}$ and $w_{a}$ take values in their entire allowed region without preference for a particular value. This comes from the narrow ranges in $z$ for each bin which limit the evolution of $f(z)$ in equation (2.2) and therefore the possible observable effects from a dynamical Dark Energy. In subsection 5.3 we do find limits for $w_{0}$ and $w_{a}$ using the full redshift range of the data.
Not all bins fit all parameters due to the fact that each bin may or may not contain a certain type of data. The only parameters fitted in all bins are $\Omega_{m}$, $H_{0}$, $w_{0}$, $w_{a}$, and the absolute magnitude of SNe Ia $M$.
5 Results
We show the fitted values of $H_{0}$ for each bin in Table 2 and illustrate them in figure 1. In all three analyses, $H_{0}$ decreases as function of redshift for $z<0.6$, which agrees with [25]. In the next two bins, the value of $H_{0}$ increases to a maximum of $\sim 78$ km $\text{s}^{-1}$ $\text{Mpc}^{-1}$ and then decreases again. Note that the scarcity and low quality of data inflate the errors at high redshifts.
[…]
Data $H_{0}\left[\frac{\text{km}}{\text{s Mpc}}\right]$
Bin 1 Bin 2 Bin 3 Bin 4 Bin 5 Bin 6 Bin 7 Bin 8
Base+PantheonPlus $73.32^{+3.34}_{-2.92}$ $67.84^{+5.37}_{-5.20}$ $66.08^{+5.26}_{-5.11}$ $56.24^{+6.91}_{-6.36}$ $67.19^{+8.92}_{-7.82}$ $78.86^{+10.30}_{-8.62}$ $64.32^{+10.24}_{-7.91}$ $64.16^{+15.82}_{-12.51}$
Base+Union3 $73.38^{+3.12}_{-2.82}$ $63.35^{+5.58}_{-5.03}$ $66.88^{+5.26}_{-5.17}$ $59.20^{+7.26}_{-6.53}$ $65.12^{+8.86}_{-7.63}$ $79.14^{+10.38}_{-9.13}$ $64.03^{+9.69}_{-8.14}$ $66.11^{+14.32}_{-11.77}$
Base+DES $73.43\pm 2.97$ $65.67^{+5.97}_{-5.52}$ $63.04^{+5.41}_{-5.11}$ $61.78^{+6.45}_{-5.99}$ $66.36^{+8.37}_{-8.09}$ $77.04^{+9.79}_{-9.02}$
Table 2: Numerical values for the best-fit $H_{0}$ in each bin and for each data set. To see the redshift range of each of the bins refer to table 1. This data is plotted in fig. 1.
In figure 1 we see that the estimations in $H_{0}$ seem to depend on the redshift. In order to quantify how much this measurements deviate from a constant, we introduce different parameterizations
$\widetilde{H}_{0}(\bar{z},\theta_{k})\,,$ (5.1)
that are functions of the redshift of the measurements $\bar{z}$ and some parameters $\theta_{k}$ for $k=1,2,\cdots$ which allow $\widetilde{H}_{0}$ to deviate from a constant. Assuming statistical independence we compute the total posterior as
$\text{P}_{\text{tot}}(\theta_{k})=\prod_{i=1}^{i_{max}}\text{P}_{i}\left(\widetilde{H}_{0}(\bar{z}_{i},\theta_{k})\right)\,,$ (5.2)
where the posteriors $P_{i}$ result from the fits of each of the bins. With this, we fit the parameters $\theta_{k}$ and estimate how much $\widetilde{H}_{0}$ deviates from a constant.
We evaluate how statistically significant is the deviation from a constant $H_{0}$ by using three different methods. First, we use the Bayesian evidence given by [58]
[…]
where $\Pi$ is the prior and $\mathscr{L}$ the Likelihood of the parameters $\theta_{k}$. We compute the integral using the nested integration algorithm Polychord [59, 60]. Here a larger value means a model with more probability. We report the Bayes factor given by
$B=\frac{E_{\text{var}}}{E_{\text{cons}}}$ (5.4)
and use it to determine if either the variable model or the constant is favored by the data.
Secondly, the Akaike Information Criterion (AIC) defined as [61, 62]
$\text{AIC}=-2\log\mathscr{L}_{\text{max}}+2k,$ (5.5)
where $\mathscr{L}_{\text{max}}$ is the maximum likelihood that can be obtained within the model and $k$ the number of free parameters. Between the two models, the variable and the constant, the one with the smaller AIC value is preferred by the data. This criterion favors the larger $\mathscr{L}_{\text{max}}$ but penalizes too many extra parameters. In the next subsections we report $\Delta\text{AIC}=\text{AIC}_{\text{var}}-\text{AIC}_{\text{cons}}$.
Thirdly, we determine the number of standard deviations ($\sigma$) away from a constant. For parameterizations with a single extra parameter we only divide the distance to a constant by the standard deviation of the parameter. For models with more parameters, we compute the $p$-value of the variable model and, assuming a Gaussian distribution, we determine the number of standard deviations associated with that $p$-value. This is the least rigorous of our tests, but serves to communicate the relevance of our results as the cosmological community is familiar with the concept that a few $\sigma$’s can be the result of statistical fluctuations and not a physical effect.
5.1 Quadratic parameterization
The first 6 bins in figure 1 seem to indicate a quadratic behaviour, therefore we propose a quadratic parameterization of the form
$\widetilde{H}_{0}(z)=\hat{H}_{0}+mz+bz^{2},$ (5.6)
with parameters $\hat{H}_{0}$, $m$ and $b$. For the fit, we exclude the results for bins 7 and 8 in order to avoid biases caused by the abrupt drop in the value of $H_{0}$ in bin 7, and the large error bars in bin 8.
[…]
From the table 3 we can see that the parameters $m$ and $b$ deviate from zero, favouring a non constant $H_{0}$ model. The values of the Bayes factor and the $\Delta$AIC however, show that this parametrization improves very modestly the fit to the data given by the constant $H_{0}$. Only accounting for a barely worth mentioning result according to the Jeffreys’ scale [63] of the Bayes factor. And according to the $\Delta$AIC scale in [62] corresponding to no preference between the models.
It is evident that the low values of $H_{0}$ in the last two bins can lead to a fitting which is compatible with a constant $H_{0}$ ($m=0$ and $b=0$). This can be considered a limitation of the quadratic parameterization, which asks for $H_{0}$ to grow up to infinity at large redshifts. As this parameterization is purely phenomenological, we change it for one that won’t present this behaviour in the next subsection.
Data $\hat{H}_{0}\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $m\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $b\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $\sigma$ $\Delta$AIC $B$
Base+PantheonPlus $75.18^{+3.76}_{-3.42}$ $-65.96^{+26.31}_{-26.69}$ $72.27^{+31.63}_{-30.81}$ 1.5 -1.67 1.59
Base+Union3 $75.06^{+3.79}_{-3.47}$ $-63.98^{+29.04}_{-28.90}$ $70.32^{+34.48}_{-33.89}$ 1.3 -0.70 1.30
Base+DES $74.64^{+3.62}_{-3.36}$ $-64.89^{+26.68}_{-26.49}$ $71.83^{+32.00}_{-32.08}$ 1.5 -1.52 1.72
Table 3: Constraints on the parameters $\hat{H}_{0}$, $m$ and $b$ in the quadratic parameterization (5.6) for the first 6 bins. We compared the model to a constant $H_{0}$ and determine the Bayes factor $B$, $\Delta$AIC and $\sigma$ deviation. While the $\sigma$ values favors a dynamic $H_{0}$, the Bayes factor indicates only a barely worth mentioning preference for the quadratic model, and the $\Delta$AIC shows no preference between them, as determined using the scale provided by [62].
[…]
5.2 Fourier parameterization
If we consider all the bins in figure 1, we see a possible oscillatory behavior. Therefore, we follow ref. [64] by proposing a parameterization in terms of the Fourier series
$\widetilde{H}_{0}(z,b_{k},c_{k})=\hat{H}_{0}+\sum_{k}\left[b_{k}\sin(2k\pi a)+c_{k}\cos(2k\pi a)\right],$ (5.7)
where $a=1/(1+z)$ is the scale factor. After fitting different cuts of the series, we have obtained that the second term in the cosine series $c_{2}$ follows the oscillatory trend particularly well. In order to simplify the parameterization we stay with
$\widetilde{H}_{0}(z)=\hat{H}_{0}+A\cos(4\pi a).$ (5.8)
with only the two parameters $\hat{H}_{0}$ and $A$.
We now fit all the available bins showing the results in table 4. Figure LABEL:fig:ajustes_modelo_cos2x displays the fit for the 8 bins in the Base+PantheonPlus and Base+Union3 samples, and for only 6 bins in the Base+DES sample, remembering this data set has a reduced redshift range.
Again, table 4 includes the $\sigma$, $\Delta$AIC and $B$ values to see the robustness of this model against a constant $H_{0}$. The parameter $A$ is different from zero up to $\sim 2.3$ standard deviations, suggesting a preference for the non constant model. However, the scale in [62] states that the difference in AIC values is insufficient to favor any of the models. Table 4 also contains the results in each data set for only 6 bins, confirming a general improvement over the quadratic parameterization. On Jeffreys’ scale [63], the Bayes factors in table 4 account for a barely worth mentioning value, with the exception of the Base+DES sample, where we found a substantial preference for the non constant model. It is interesting to note that the latest two bins in Base+PantheonPlus and Base+Union3 don’t change too much the results in each data set compared to 6 bins.
Data $\hat{H}_{0}\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $A\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $\sigma$ $\Delta$AIC $B$
6 bins
Base+PantheonPlus $67.15^{+2.40}_{-2.35}$ $5.70^{+2.82}_{-2.88}$ 2 $-1.90$ 2.11
Base+Union3 $67.10^{+2.34}_{-2.47}$ $5.35^{+3.08}_{-2.94}$ $1.8$ $-1.22$ 1.90
Base+DES $66.15^{+2.39}_{-2.19}$ $6.53^{+2.88}_{-2.85}$ $2.3$ $-2.90$ 3.87
8 bins
Base+PantheonPlus $66.63^{+2.39}_{-2.27}$ $5.59^{+2.76}_{-2.78}$ 2 $-2.03$ 2.21
Base+Union3 $66.68^{+2.25}_{-2.34}$ $5.24^{+2.97}_{-2.91}$ $1.8$ $-1.26$ 1.54
Table 4: Constraints on the parameters $\hat{H}_{0}$ and $A$ in the Fourier parameterization (5.8). Remember that Base+PantheonPlus and Base+Union3 reach up to $z=2.3$ while Base+DES only up to $z=1.13$. We compared the model to a constant $H_{0}$ and determine the Bayes factor $B$, $\Delta$AIC and $\sigma$ deviation. We see that the deviation from a constant is larger in $\sigma$’s than for the quadratic model. The $\Delta$AIC is slightly better for the Fourier model than for a constant, although from the scale in [62] there is no preference to either model. The Bayes factors values account for a barely worth mentioning value according to the Jeffreys’ scale [63], with the exception of the Base+DES data, which shows a substantial preference for the non constant model.
5.3 Unbinned results
[…]
Data ${H}_{0}\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $\Omega_{m}$ $w_{o}$ $w_{a}$ $r_{d}$ [Mpc]
Base+PantheonPlus $71.08^{+2.15}_{-2.07}$ $0.31\pm 0.02$ $-0.88^{+0.08}_{-0.07}$ $-0.45^{+0.58}_{-0.61}$ $140.95^{+4.29}_{-4.11}$
Base+Union3 $70.39^{+2.16}_{-2.11}$ $0.33\pm 0.02$ $-0.68\pm 0.13$ $-1.35^{+0.74}_{-0.78}$ $139.58^{+4.31}_{-4.13}$
Base+DES $70.63^{+2.13}_{-2.08}$ $0.33\pm 0.02$ $-0.75\pm 0.09$ $-1.17^{+0.64}_{-0.66}$ $140.13^{+4.26}_{-4.20}$
Table 5: Best-fit values on the parameters $H_{0}$, $\Omega_{m}$, $w_{0}$, $w_{a}$ and $r_{d}$ of the megamasers+CC+DESI+SNe dataset, without binning and over the entire redshift range $z<2.33$. Without the CC and megamasers data, $H_{0}$, $r_{d}$ and $M$ cannot be determined. We see that the constraints on $w_{0}$ and $w_{a}$ resemble those in [35], for DESI + SNe data. A cosmological constant ($w_{0}=-1$ and $w_{a}=0$) is excluded by $\sim 2\sigma$ when the SNe data are either Union3 or DES.
6 Summary and conclusions
In this work, we have studied the evolution in the value of $H_{0}$ with the redshift of the data used to determine it, finding a dynamic evolution in $H_{0}$. For that we used the latest dataset of CC, megamaser, SNe Ia and BAO observations from the Year 1 data release of DESI. The results show that the Hubble tension, referring to a phenomenologically non-constant value in $H_{0}$, cannot be resolved by the new DESI data together with a dynamical DE.
The SNe observations are chosen from one of the three most recent samples of Pantheon+, Union3 and DES. Pantheon+ and Union 3 share about 1360 SNe but differ in their analysis and treatment over systematics. DES has 1635 new SNe but share 194 low redshift SNe with Pantheon+. Our three data samples share the megamasers+CC+DESI data but differ in the SNe data. Base+PantheonPlus and Base+Union3 formed a total of 8 bins, and 6 bins for Base+DES.
[…]
The results show a dynamical trend in $H_{0}$ with a statistical significance up to $1.5\sigma$, for the quadratic parameterization (5.6) and up to $2.3\sigma$ for the Fourier parameterization (5.8), indicating a preference for a non constant model. However, the values in $\Delta$AIC and the Bayes factor $B$, which penalize the extra parameters in the dynamical models, do not provide substantial support for the quadratic function against a constant value. From a Bayesian point of view, both the quadratic function and a constant $H_{0}$ fit the bins results equally well. Similar conclusions are obtained for the Fourier parameterization, except for the results of Base+DES, which have the highest $\Delta$AIC and $B$ values showing a substantial preference for the dynamical model, according to the Jeffrey’s scale.
[…]
We highlight that the three data sets, which use different SNe samples, show similar behaviors even though each sample differs in its content and methodology. Because of this, we can infer that this trend is not caused by uncountable systematic errors in the SNe parameters and it is more of an intrinsic behavior guided by the data.
We have ignored any calibration on either $r_{d}$ or M. However, if we adopt them, we find that the bins values of $H_{0}$ are completely consistent with a constant. This would require to use data outside of the redshift bin, either CMB data to determine $r_{d}$ or distance leader data to determine M. As fixing $r_{d}$ or M is equivalent to fixing $H_{0}$ from the beginning, which prevents the data from speaking for itself. It is relevant to highlight that any solution to the Hubble tension that only modifies the calibration of $r_{d}$ or M won’t solve the tension presented here. In the unbinned case, in subsection 5.3 we obtained a $r_{d}$ which is $1.5\sigma$ smaller than the one measured by Planck.
Of course, the results for $H_{0}$ may vary if one adopts a different binning strategy but we expect that the significance of the trend do not change significantly. This evolution in the Hubble constant might offer a new perspective on the Hubble tension. Since we’re using the $w_{0}w_{a}$ model with very broad priors, it seems that a dynamical DE can’t remove this trend. Using a cosmological constant as the Dark Energy component only reduces the errorbars of the estimation of $H_{0}$ due to the reduced parameter space. This may increase the significance of the non constant models.
Some possible solutions would be either a hidden systematic error or some other assumption in the model. This could even be the reason why SH0ES and Planck have different values for $H_{0}$ given that they use data with different redshifts, but this needs further study, maybe considering $H_{0}$ not as an integration constant, but as a quantity dependent on the observed region of the Universe. Unfortunately, directly corroborating that the Hubble parameter at present time has different values depending on the region of the Universe is difficult, as due to the finiteness of the speed of light, we observe those regions in the past. However, nothing prevents us from thinking that just as there are perturbations to matter, these bring with them anisotropies in the expansion rate of the Universe with $H_{0}$ being more of an average value and the local departure of the cosmological principle [65, 66, 67, 68, 69, 70].
Our binning strategy splits the data used, reducing the available data in each bin. This, combined with the decreasing quality in the data, produces large error bars in the inferred parameters which get worse at large $z$’s. Therefore the quadratic and Fourier fittings present large uncertainties. We do not rule out that the constant $H_{0}$ be the correct result; we need to wait to improve the quality and number of the data to be able to reach a final conclusion.
The Hubble tension remains one of the most significant challenges in modern cosmology and astrophysics, highlighting persistent discrepancies in $H_{0}$ values derived from different observational methods and across varied redshifts. This work contributes to the present discussion, providing further evidence that the Hubble tension is not only confined to early- versus late-time measurements but also appears consistently within purely late-time measurements reflecting unresolved aspect of cosmological models and data calibration.
Acknowledgments
The authors thankfully acknowledge Sebastien Fromenteau for his useful comments in the development of this work. M. L. H. was supported by Conahcyt grant 806098.